# Nambu-Goto equation from three-dimensional gravity

### v2

@@ 1 Introduction
Our results lead to a fresh perspective on Einstein’s dilemma of whether to consider the right hand side of his eponymous equations of gravity as ugly. The usual formulation of string theory prioritizes the right hand side. It geometrizes and unifies matter as quantum vibrations of a fundamental string, and also makes semi-classical gravitational spacetimes emerge from the worldsheet. However, our results imply that the classical motion of the fundamental string is itself a consequence of spacetime dynamics. Therefore, it is natural to ask whether the spectrum and quantum dynamics of the (first quantized) fundamental string Polchinski:1998rq ; Polchinski:1998rr ; Maldacena:2000hw ; Maldacena:2000kv ; Gaberdiel:2018rqv can also be part of the quantization of pure gravity where junctions with a finite tension are included.
@@ 5 Discussion
Another fundamental question is whether we can quantize our three-dimensional gravitational solutions with junctions generalizing the approach in Kim:2015qoa ; Cotler:2018zff ; Collier:2023fwi (based on rewriting three-dimensional gravity in AdS as PSL$(2,\mathbb{R})$$\times$ PSL$(2,\mathbb{R})$ Chern-Simons theory Achucarro:1986uwr ; Witten:1988hc and further work in Krasnov:2005dm ; Scarinci:2011np ) and obtain the spectrum of first quantized bosonic string theory in background flat space Polchinski:1998rq ; Polchinski:1998rr or AdS space Maldacena:2000hw ; Maldacena:2000kv ; Gaberdiel:2018rqv within the full Hilbert space. The latter should however be bigger and include boundary gravitons that scatter off the quantum string. Quantum string theory can thus emerge directly from lower dimensional gravity together with additional degrees of freedom. The presence of rigid deformations of the Nambu-Goto equations arising from the junction conditions implies that the quantized string theory obtained from quantizing such gravitational solutions can have additional non-trivial features as well. Of course, we can generalize the junctions in the context of gauged supergravities and wonder if the classical and quantum superstring theory can emerge similarly from super-spacetime dynamics.

### v3

@@ 1 Introduction
Our results lead to a fresh perspective on Einstein’s dilemma of whether to consider the right hand side of his eponymous equations of gravity as ugly. The usual formulation of string theory prioritizes the right hand side. It geometrizes and unifies matter as quantum vibrations of a fundamental string, and also makes semi-classical gravitational spacetimes emerge from the worldsheet. However, our results imply that the classical motion of the fundamental string is itself a consequence of spacetime dynamics in three dimensions. Therefore, it is natural to ask whether the spectrum and quantum dynamics of the (first quantized) fundamental string Polchinski:1998rq ; Polchinski:1998rr ; Maldacena:2000hw ; Maldacena:2000kv ; Gaberdiel:2018rqv can also be part of quantized pure three-dimensional gravity in which junctions with a finite tension are included. [^2] Since we are discussing three-dimensional gravity, we cannot hope to recover the full spectrum of the fundamental string in critical number of dimensions.
@@ 5 Discussion
Another fundamental question is whether we can quantize our three-dimensional gravitational solutions with junctions generalizing the approach in Kim:2015qoa ; Cotler:2018zff ; Collier:2023fwi (based on rewriting three-dimensional gravity in AdS as PSL$(2,\mathbb{R})$$\times$ PSL$(2,\mathbb{R})$ Chern-Simons theory Achucarro:1986uwr ; Witten:1988hc and further work in Krasnov:2005dm ; Scarinci:2011np ) and obtain part of the spectrum of first quantized bosonic string theory in background flat space Polchinski:1998rq ; Polchinski:1998rr or AdS space Maldacena:2000hw ; Maldacena:2000kv ; Gaberdiel:2018rqv within the full Hilbert space. The latter should however also include boundary gravitons that scatter off the quantum string. Aspects of quantum string theory can thus emerge directly from lower dimensional gravity together with additional degrees of freedom. The presence of rigid deformations of the Nambu-Goto equations arising from the junction conditions implies that the quantized string theory obtained from quantizing such gravitational solutions can have other non-trivial features as well. Of course, we can generalize the junctions in the context of gauged supergravities and wonder if some aspects of classical and quantum superstring theory can emerge similarly from super-spacetime dynamics in three dimensions.